Variable Cost Method of Product Pricing

Smart Stream Inc. produces and sells cell phones. The costs of producing and selling 4,000 units of cell phones are as follows:

Variable costs: Fixed costs:
    Direct materials $ 60 per unit     Factory overhead $99,000
    Direct labor 28     Selling and admin. exp. 34,800
    Factory overhead 18
    Selling and admin. exp. 14
     Total variable cost per unit $120 per unit

Smart Stream Inc. desires a profit equal to a 15% rate of return on invested assets of $420,000.

Assume that Smart Stream Inc. uses the variable cost method of applying the cost-plus approach to product pricing.

a. Determine the variable costs and the variable cost amount per unit for the production and sale of 4,000 units of cellular phones.

Total variable cost $
Variable cost amount per unit $

b. Determine the variable cost markup percentage for cellular phones.
%

c. Determine the selling price of cellular phones. Round to the nearest cent.
$ per cellular phone
Answer #1
a
Total variable cost 480000 =120*4000
Variable cost amount per unit 120
b
Required markup 196800 =(420000*15%)+99000+34800
Divide by Total variable cost 480000
Variable cost markup percentage 41.00%
c
Variable cost amount per unit 120.00
Add: Markup per unit 49.20 =120*41%
Selling price 169.20 per cellular phone